would you please provide anwsers and show how is done thank you
1. (6±√32)÷3 simplify
2. (16±√56)÷20 simplified
3. 8+10x-3xsquard=0 factoring
4. 25xsquard=16 extracting square roots


Answer by edjones(8007)

(6±√32)÷3
(6+-sqrt(16*2))/3
(6+-4sqrt(2))/3
2+-(4/3)sqrt(2)
.
25x^2=16
5x=+-4
x=4/5, x=-4/5
.
-3x^2+10x+8=0
3x^2-10x-8=0
3*-8=-24 What 2 factors of -24 add up to -10? Answer: -12 & +2
3x^2-12x+2x-8=0
3x(x-4)+2(x-4)=0
(3x+2)(x-4)=0
x=4
3x=-2
x=-2/3
.
Ed